# LLMSteer: Improving Long-Context LLM Inference by Steering Attention on Reused Contexts

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex tasks such as question answering, summarization, and reasoning  llm (a, b, c). To enhance their reliability, LLMs are often augmented with domain-specific or user-specific knowledge that extends beyond their inherent training data Lewis et al. (2020); Jiang et al. (2023); Chen et al. (2024). However, incorporating these supplemental contexts, which can exceed thousands of tokens Jin et al. (2024); Gao et al. (2023), presents two challenges: (1) models often struggle to comprehend long context (e.g., lost-in-the-middle problem Liu et al. (2023a); Junqing et al. (2023)) and (2) processing long context incurs substantial runtime costs Liu et al. (2024); Lin et al. (2024); Zhong et al. (2024).
Since the Key-Value (KV) cache of the same context text chunks is often reused Liu et al. (2023b); Yao et al. (2024); Jin et al. (2024) multiple times, many recent systems adopt prefix cachingJin et al. (2024); Liu et al. (2023b); Qin et al. (2024), which stores the KV caches for the frequently reused contexts such that LLMs no longer need to prefill these contexts repeatedly. However, the model persists in losing track of key information from the context as its KV pairs remain unchanged. So, is there a way to simultaneously achieve high efficiency and high quality without fine-tuning models?
@@
We implemented LLMSteer on top of LLaMA-3.1-8b-Instruct Dubey et al. (2024), and compared LLMSteer with LLaMA-3.1-8b-Instruct, LLaMA-3.1-70b-Instruct Dubey et al. (2024) and the state-of-the-art attention steering baseline. LLMSteer demonstrates that by steering tokens that are consistently important across the model’s understandings, it not only achieves a significant increase in F1 scores (e.g., 72.9 -> 82.0), but also delivers up to 4.8× faster compared to the baselines. LLMSteer not only reduces runtime costs but also improves generation quality. This paper is the first effort to (1) improve model generation quality without fine-tuning and (2) do so in a way that is compatible with prefix caching.
@@ 2 Post-hoc attention steering
Attention steering didn’t come out of nowhere. Methods like PASTA Zhang et al. (2023) and AutoPASTA Anonymous (2024) introduce techniques to boost the generation quality of LLMs by guiding their attention to specific parts of the input text. In PASTA, users manually specify tokens in the input and steer the LLM’s attention toward these highlighted tokens through reweighting. However, annotating relevant input spans by humans is not always feasible, especially for lengthy contexts and context-specific tasks.
@@
First, the iterative prompting approach necessitates two LLM calls, which increases runtime latency. Second, AutoPASTA is incompatible with prefix caching, a technique in which KV caches of the frequently used text chunks are stored and reused such that the prefill of the reused text chunks can be skipped. For example, Llama-8B Dubey et al. (2024) requires 2.04 seconds to process a 5000-token context using 2 A40 GPUs. However, if the KV cache of the text chunk is precomputed and stored, only the final user query needs to be prefilled, which only takes 0.039 seconds on the same hardware. Although prefix caching can still be applied to the initial call in AutoPASTA, the second call cannot benefit from prefix caching in that the attention of the context is steered in a query-dependent way. Consequently, the KV cache of the context will be altered and can no longer be reused across different queries.
@@ 4.1 Setup
Datasets:  Our evaluation covers three datasets: SQuAD Rajpurkar et al. (2016), TriviaQA Joshi et al. (2017) and GSM8K Cobbe et al. (2021). SQuAD is a reading comprehension dataset with questions based on Wikipedia articles. TriviaQA is a large reading comprehension dataset, requiring reasoning across multiple sentences. GSM8K is a collection of diverse grade school math word problems designed for multi-step reasoning. For each dataset, we randomly sample 100 test cases. We use F1-score as the quality metric and request delay (in seconds) as the efficiency metric across all datasets.
@@
Implementation:  We build LLMSteer and all the baselines on top of Huggingface Transformers Wolf et al. (2020). We implement AutoPASTA based on their paper as the code is not public yet. All experiments are conducted on two NVIDIA A40 GPUs.
@@ 5 Limitations and Future work
We plan to extend to longer context lengths (e.g., >10k tokens) to fully explore LLMSteer’s capabilities and to test it on models beyond Llama-8B to assess generalizability. We will also conducting an ablation study to quantify the contributions of the steering mechanism versus contextual re-reading. We will explore the limitations and capabilities of our method compared to traditional fine-tuning, especially considering context window length and model reasoning capabilities. Additionally, PagedAttention Kwon et al. (2023) and FlashAttention Dao et al. (2022) could enhance efficiency. Investigating more fine-grained steering at the granularity of individual token pairs, beyond our current token-level attention upweighting, may further improve generation quality and will be explored in future work.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex tasks such as question answering, summarization, and reasoning ( llm (a, b, c)). To enhance their reliability, LLMs are often augmented with domain-specific or user-specific knowledge that extends beyond their inherent training data ( Lewis et al. (2020); Jiang et al. (2023); Chen et al. (2024)). However, incorporating these supplemental contexts, which can exceed thousands of tokens ( Jin et al. (2024); Gao et al. (2023)), presents two challenges: (1) models often struggle to comprehend long context (e.g., lost-in-the-middle problem ( Liu et al. (2023a); Junqing et al. (2023))) and (2) processing long context incurs substantial runtime costs (Liu et al. (2024); Lin et al. (2024); Zhong et al. (2024)).
Since the Key-Value (KV) cache of the same context text chunks is often reused multiple times (Liu et al. (2023b); Yao et al. (2024); Jin et al. (2024)), many recent systems adopt prefix caching (Jin et al. (2024); Liu et al. (2023b); Qin et al. (2024)), which stores the KV caches for the frequently reused contexts such that LLMs no longer need to prefill these contexts repeatedly. However, the model persists in losing track of key information from the context as its KV pairs remain unchanged. So, is there a way to simultaneously achieve high efficiency and high quality without fine-tuning models?
@@
We implemented LLMSteer on top of LLaMA-3.1-8b-Instruct ( Dubey et al. (2024)), and compared LLMSteer with LLaMA-3.1-8b-Instruct, LLaMA-3.1-70b-Instruct ( Dubey et al. (2024)) and the state-of-the-art attention steering baseline. LLMSteer demonstrates that by steering tokens that are consistently important across the model’s understandings, it not only achieves a significant increase in F1 scores (e.g., 72.9 -> 82.0), but also delivers up to 4.8× faster compared to the baselines. LLMSteer not only reduces runtime costs but also improves generation quality. This paper is the first effort to (1) improve model generation quality without fine-tuning and (2) do so in a way that is compatible with prefix caching.
@@ 2 Post-hoc attention steering
Attention steering didn’t come out of nowhere. Methods like PASTA ( Zhang et al. (2023)) and AutoPASTA ( Anonymous (2024)) introduce techniques to boost the generation quality of LLMs by guiding their attention to specific parts of the input text. In PASTA, users manually specify tokens in the input and steer the LLM’s attention toward these highlighted tokens through reweighting. However, annotating relevant input spans by humans is not always feasible, especially for lengthy contexts and context-specific tasks.
@@
First, the iterative prompting approach necessitates two LLM calls, which increases runtime latency. Second, AutoPASTA is incompatible with prefix caching, a technique in which KV caches of the frequently used text chunks are stored and reused such that the prefill of the reused text chunks can be skipped. For example, Llama-8B ( Dubey et al. (2024)) requires 2.04 seconds to process a 5000-token context using 2 A40 GPUs. However, if the KV cache of the text chunk is precomputed and stored, only the final user query needs to be prefilled, which only takes 0.039 seconds on the same hardware. Although prefix caching can still be applied to the initial call in AutoPASTA, the second call cannot benefit from prefix caching in that the attention of the context is steered in a query-dependent way. Consequently, the KV cache of the context will be altered and can no longer be reused across different queries.
@@ 4.1 Setup
Datasets:  Our evaluation covers three datasets: SQuAD ( Rajpurkar et al. (2016)), TriviaQA ( Joshi et al. (2017)) and GSM8K ( Cobbe et al. (2021)). SQuAD is a reading comprehension dataset with questions based on Wikipedia articles. TriviaQA is a large reading comprehension dataset, requiring reasoning across multiple sentences. GSM8K is a collection of diverse grade school math word problems designed for multi-step reasoning. For each dataset, we randomly sample 100 test cases. We use F1-score as the quality metric and request delay (in seconds) as the efficiency metric across all datasets.
@@
Implementation:  We build LLMSteer and all the baselines on top of Huggingface Transformers ( Wolf et al. (2020)). We implement AutoPASTA based on their paper as the code is not public yet. All experiments are conducted on two NVIDIA A40 GPUs.
@@ 5 Limitations and Future work
We plan to extend to longer context lengths (e.g., >10k tokens) to fully explore LLMSteer’s capabilities and to test it on models beyond Llama-8B to assess generalizability. We will also conducting an ablation study to quantify the contributions of the steering mechanism versus contextual re-reading. We will explore the limitations and capabilities of our method compared to traditional fine-tuning, especially considering context window length and model reasoning capabilities. Additionally, PagedAttention (Kwon et al. (2023)) and FlashAttention (Dao et al. (2022)) could enhance efficiency. Investigating more fine-grained steering at the granularity of individual token pairs, beyond our current token-level attention upweighting, may further improve generation quality and will be explored in future work.